OpenStudy (anonymous):

Find the distance between the points (13, 20) and (18, 8)

jimthompson5910 (jim_thompson5910):

Use the distance formula to get \[\Large d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] \[\Large d = \sqrt{(13-18)^2+(20-8)^2}\] \[\Large d = \sqrt{(-5)^2+(12)^2}\] \[\Large d = ???\]

OpenStudy (♪chibiterasu):

Use the distance formula: \[(x _{1} - x _{2}) + (y _{1} - y _{2})\]

OpenStudy (♪chibiterasu):

pff, yeah that.

OpenStudy (♪chibiterasu):

I don't think I wrote it right anyway lol.